Mitch and Bill are both age 75. When Mitch was 22 years​ old, he began depositing ​$1300 per year into a savings account. He made deposits for the first 10​ years, at which point he was forced to stop making deposits.​ However, he left his money in the​ account, where it continued to earn interest for the next 43 years. Bill​ didn't start saving until he was 46 years​ old, but for the next 29 years he made annual deposits of ​$1300. Assume that both accounts earned an average annual return of 5​% ​(compounded once a​ year). Complete parts​ (a) through​ (d) below.
a. How much money does Mitch have in his account at age​ 75?
At age​ 75, Mitch has ​$
nothing in his account.
​(Round to the nearest cent as​ needed.)

Answer by Boreal(15235)

Mitch P=1300((1.05)^10-1)/.05=$16,351.26
for the next 43 years, 16351.26(1.05)^43=$133,257.33
Bill P=1300((1.05)^29-1)/.05=$81,019.53